OpenStudy (anonymous):

I don't get this question someone please explain Which values will make the equation true? −3/5=? Drag each value to the table to show whether it makes the equation true or false. Makes the equations true: Makes them false: -3/5 3/5 3/-5 -( -3/-5) -(-3/5) -(3/5)

OpenStudy (anonymous):

@AlexandervonHumboldt2 @jim_thompson5910

OpenStudy (aravindg):

The question doesn't look right.

OpenStudy (♪chibiterasu):

Which values are the same as -3/5

OpenStudy (anonymous):

Ok so how do I do that?

OpenStudy (anonymous):

For a decimal I got .6 , does that help?

OpenStudy (♪chibiterasu):

-3/5 3/5 3/-5 -(-3/-5) -(3/5) Which one of these obviously equals -3/5 that doesnt help we just need to take into account the signs

OpenStudy (♪chibiterasu):

Thats one way to do it but this way is much simpler

OpenStudy (anonymous):

What so -3/5= -3/5??

OpenStudy (♪chibiterasu):

yes

OpenStudy (♪chibiterasu):

Now the next one does -3/5 = 3/5?

OpenStudy (anonymous):

wait hold on, the fraction says -3/5 for whole and the 2nd fraction literally said -3/5 as in negative 3 over 5

OpenStudy (♪chibiterasu):

what

OpenStudy (anonymous):

exactly

OpenStudy (♪chibiterasu):

i dont understand what you're saying

OpenStudy (anonymous):

um i'll try to explain it more easily

OpenStudy (♪chibiterasu):

can you just screenshot it?

OpenStudy (anonymous):

I have a toshiba laptop idk how to

OpenStudy (♪chibiterasu):

Ok I get what you mean, but when one of the values is negative it makes the entire fraction negative so its the same thing

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So are they both true?

OpenStudy (♪chibiterasu):

So when both of them are negative it cancels out and it becomes positive

OpenStudy (♪chibiterasu):

-3/5 = -3/5 so yes

OpenStudy (♪chibiterasu):

but does -3/5 = 3/5?

OpenStudy (anonymous):

Um no?

OpenStudy (♪chibiterasu):

yeah. so you would drag that into false

OpenStudy (anonymous):

do I just drag the rest into false other than-3/5 and -3/5?

OpenStudy (anonymous):

i mean

OpenStudy (♪chibiterasu):

no were not done now 3/-5 remember what I said about one value being negative?

OpenStudy (anonymous):

just -3/5

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

So true?

OpenStudy (♪chibiterasu):

we're not done we can still keep going. signs can be tricky

OpenStudy (♪chibiterasu):

yes, 3/-5 would be true. -3/5 = 3/-5 because both will just end up being -(3/5) when simplified

OpenStudy (♪chibiterasu):

do you kinda get it now?

OpenStudy (anonymous):

Yeah, thanks so much for helping with explaining the problem

OpenStudy (♪chibiterasu):

were not done!

OpenStudy (anonymous):

o

OpenStudy (♪chibiterasu):

-( -3/-5) what do you think this means?

OpenStudy (anonymous):

Um not sure, is it true also? or is it false

OpenStudy (anonymous):

i think false

OpenStudy (♪chibiterasu):

First lets look inside the parenthesis remember that a - sign is just reversing signs. It will turn - into + and it will turn + into -. in the parenthesis, both values are negative. do you remember waht i said about that?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so It is true?

OpenStudy (♪chibiterasu):

yeah. -3/-5 turns into 3/5 so now we have -(3/5) which does equal -3/5

OpenStudy (anonymous):

Ah I see

OpenStudy (♪chibiterasu):

-(-3/5) what about this one?

OpenStudy (anonymous):

True? Because it changes -3 into 3 so -(3/5)?

OpenStudy (♪chibiterasu):

it does change -3 into 3 but after you convert you can get rid of the parenthesis so it would just be 3/5

OpenStudy (anonymous):

oh then what happened to the negative sign by the parenthesis?

OpenStudy (♪chibiterasu):

You don't need it anymore after you convert signs, so we get rid of it when we convert -3 into 3 :)

OpenStudy (♪chibiterasu):

It's kinda like using a power-up. Once you use it, it vanishes or loses its purpose and isn't necessary anymore.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

so it's false then?

OpenStudy (♪chibiterasu):

yeap

OpenStudy (♪chibiterasu):

last one -(3/5) i think this one is pretty obvious lol

OpenStudy (anonymous):

True haha

OpenStudy (anonymous):

wait, right?

OpenStudy (♪chibiterasu):

yeah, it's right xD

OpenStudy (anonymous):

Oh ok lol, thanks again for helping me! I learned a lot about fractions during this process :D

OpenStudy (♪chibiterasu):

so just to make sure, what did you put in for true and false?

OpenStudy (♪chibiterasu):

im glad :)

OpenStudy (anonymous):

Oh I put

OpenStudy (anonymous):

um Makes the equation true:-(-3/-5) , -(3/5) , -(-3/5) , 3/-5 , -3/5 Is that right?

OpenStudy (♪chibiterasu):

-(-3/5) is false. It changes the negative to a positive :)

OpenStudy (♪chibiterasu):

the rest is fine, good job

OpenStudy (anonymous):

Ok thanks!!! You helped a lot

OpenStudy (♪chibiterasu):

np